can u please help
If the ratio of two supplementary angles is 7 to 8, calculate the measure of the larger of the two angles.

Found 2 solutions by rfer, RAY100:
Answer by rfer(12678)

small34 angle=7x=84
larger angle=8x=96
7x+8x=180
15x=180
x=12
7*12=84
8*12=96

Answer by RAY100(1637)

Supplemental means that the sum of the 2 angles is 180 degrees
.
If one angle is "x" ,,,the other angle is "180-x"
.
ratio = x/(180-x) = 7/8
.
cross multiply
.
8x = 7*180 - 7x
.
15x = 1260
.
x= 84
.
(180-x) = 180-84 = 96
.
check,,ratio 84/96 = 7/8,,,,ok
,,,,,96 - 84 = 180,,,ok